# State Combinatorial Generalization In Decision Making With Conditional Diffusion Models

### v1

@@ A.1 Generalization in RL
We prove Corollary 5.1 by construction. Suppose $\mathcal{M}$ is a $d$-dimensional linear manifold and ${\mathbf{S}}\subset\mathbb{R}^{k}$, then we note that both $\mathcal{M}$ and ${\mathbf{Z}}$ are affine isomorphic to $\mathbb{R}^{d}$. Therefore, with necessary shifting, we can have ${\bm{z}}^{\prime}={\bm{z}}+{\bm{y}}$ where ${\bm{y}}\in{\mathbf{Z}}$. Let ${\bm{v}}$ be the corresponding vector of ${\bm{y}}$ in $\mathcal{M}$.
Now let’s perform DDIM inversion on a training sample ${\bm{s}}$ to obtain the SDE trajectory $\{{\bm{s}}_{t}\}$. Let $\gamma_{t}$ be the angle between $\epsilon_{\theta}({\bm{s}}_{t},t)$ and $\mathcal{M}$, there exist a set of vectors ${\bm{v}}_{t}$ such that ${\bm{v}}=\sum_{t}sin(\gamma_{t})\sigma_{t}{\bm{v}}_{t}$ and ${\bm{v}}_{t}$ perpendicular to $\epsilon_{\theta}({\bm{s}}_{t},t)$. Then by construction and with necessary shifting, the trajectory ${\bm{s}}_{t}^{\prime}={\bm{s}}_{t}+{\bm{v}}_{t}$ is a valid diffusion denoising trajectory (with ${\bm{v}}_{t}$ acting as the “random” vector $\mathbf{\epsilon}_{t}$ sampled at each time step). This trajectory will yield ${\bm{s}}^{\prime}$ as the final state with non-zero probability because each intermediate Gaussian distribution $p_{\theta}({\bm{s}}_{t-1}|{\bm{s}}_{t})=\mathcal{N}({\bm{s}}_{t-1};\bm{\mu_{\theta}}({\bm{s}}_{t},t),\bm{\Sigma_{\theta}}({\bm{s}}_{t},t))$ is defined on the entire ambient space. By Stanczuk et al. (2022) we know that $\gamma_{t}\to\pi/2$ as $t\to 0$, therefore ${\bm{v}}$ can be a non-zero vector. Hence, there exists a ${\bm{v}}\in\mathcal{M}$ such that the sampling probability of ${\bm{s}}^{\prime}={\bm{s}}+{\bm{v}}$ from diffusion model $p_{\theta}$ is non-zero. ∎
While we have proven non-zero probability above, one can easily spot that, the probability can become extremely close to zero if the OOC sample is very far away from all training examples due to the intermediate Gaussian distributions. This corresponds to generalizing to the out-of-distribution samples with unseen base elements (the gray area in Figure 1). One can mitigate this problem by increasing the coverage of the support of the training space, which is also a common method in traditional RL to mitigate the problem of generalization to unseen base elements. Applying other post-training sampling techniques like repainting (Lugmayr et al., 2022) can also allow extra Langevin steps when $t$ is large.
Appendix C Additional Visualization of Value Function of PPO
We include here the value prediction of PPO for in-distribution and OOC states to demonstrate that OOC states are also a problem for online methods.
Figure 9: Value prediction of PPO and actual return-to-go (RTG) in unsupported states in Roundabout environment
Appendix D Experiment Details
D.1 Hardware and Platform
@@ D.4 Trajectory formulation
D.5 Pseudo-code
Pseudo-code for planning with conditional diffusion model is shown below in Algorithm 1.
Input: Diffusion model $\bm{\epsilon_{\theta}}$, compositional elements extractor $r$, learnable embedding function $h$, classifier-free guidance scale $\lambda$, state dimentionality $d_{S}$, initial observation $\bm{o}$, environment simulator $env$
while not done do
     Initialize $\bm{{\tau}_{t}}\sim\mathcal{N}(0,I)$
     $\bm{c}\leftarrow r(\bm{o})$triangleright Extract observed compositional information
     $\bm{z}\leftarrow h(\bm{c})$triangleright Obtain element embedding
     for $t\leftarrow T,...1$ do
         $\bm{{\tau}}_{t}[:d_{S},0]\leftarrow\sqrt{\bar{\alpha}_{t}}\bm{o}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},\text{ where }\bm{\epsilon}\sim\mathcal{N}(0,I)$triangleright Replace the first observed
                                                                                      state with noised $\bm{o}$
         $\bm{\widetilde{\epsilon_{t}}}=(1+\lambda)\bm{\epsilon_{\theta}}({\bm{s}}_{t},\bm{z},t)-\lambda\bm{\epsilon_{\theta}}({\bm{s}}_{t},t)$triangleright Classifier free guidance
         $\bm{{\tau}_{t-1}}=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{{\tau}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\widetilde{\epsilon_{t}}}\right)+\sigma_{t}\bm{\epsilon_{t}},\text{ where }\bm{\epsilon_{t}}\sim\mathcal{N}(0,I)$
     end for
     $\bm{a}\leftarrow\bm{{\tau}_{0}}[d_{S}:,0]$triangleright Extract action
     $\bm{o}\leftarrow env.step(\bm{a})$
end while
D.6 Maze2D
D.6.1 Extra Results
@@
D.6.2 Experiment Details
We followed the setup used in Janner et al. (2022). The hyperparameters shared for large and medium mazes are shown below in Table 1. Large maze use a planning horizon of 384 and medium maze use a planning horizon of 256. Conditioning is passed through a positional embedding layer first to map each dimension of the waypoint $(x,y,v_{x},v_{y})$ to a higher dimension of 21 and concatenate them to form a vector of size $(1,21*4)$. Three waypoints are then stacked together to form a matrix of size $(3,21*4)$ and passed into the cross-attention layer. In our experiment, directly using the waypoints as conditioning was unsuccessful.
@@ D.7.3 Dataset
D.8 Setup for Roundabout Environment
[…]
D.8.1 Experiment Details

### v2

@@ A.1 Generalization in RL
Because $g$ is block-wise bi-Lipschitz, $g$ is also injective and differentiable almost everywhere by Rademacher’s theorem. Then using the exact density transform, for almost everywhere $\bm{y}$}, we can have ${\bm{x}}^{*}=g^{-1}({\bm{y}})$ and
$p_{Y}({\bm{y}})=\frac{p_{X}({\bm{x}}^{*})}{|\operatorname{det}Dg({\bm{x}}^{*})|}$ (8)
Equation 8 entails that in order to obtain a lower bound for $p_{Y}({\bm{y}})$, we should obtain a lower bound for $p_{X}({\bm{x}}^{*})$ and an upper bound for $|\operatorname{det}\nabla_{{\bm{x}}^{*}}g({\bm{x}}^{*})|$.
From Equation 6 we can directly obtain
$p_{X}({\bm{x}}^{*})\geq c(2\pi)^{-d/2}(\sigma^{\mathcal{M}})^{-m}(\sigma^{\perp})^{m-d}\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{x}}^{*}-\mu)\|^{2}}{2(\sigma^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{x}}^{*}-\mu)\|^{2}}{2(\sigma^{\perp})^{2}}\right)}$ (9)
By the Liptschtiz condition we also know that
$A^{\mathcal{M}}\|W_{\mathcal{M}}({\bm{x}}^{*}-\mu)\|$ $\leq\|W_{\mathcal{M}}(g({\bm{x}}^{*})-g(\mu))\|\Rightarrow\|W_{\mathcal{M}}({\bm{x}}^{*}-\mu)\|\leq\frac{\|W_{\mathcal{M}}({\bm{y}}-g(\mu))\|}{A^{\mathcal{M}}}$ (10)
$A^{\perp}\|W_{\perp}({\bm{x}}^{*}-\mu)\|$ $\leq\|W_{\perp}(g({\bm{x}}^{*})-g(\mu))\|\Rightarrow\|W_{\perp}({\bm{x}}^{*}-\mu)\|\leq\frac{\|W_{\perp}({\bm{y}}-g(\mu))\|}{A^{\perp}}$
So
$p_{X}({\bm{x}}^{*})\geq c(2\pi)^{-d/2}(\sigma^{\mathcal{M}})^{-m}(\sigma^{\perp})^{m-d}\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{y}}-g(\mu))\|^{2}}{2(\sigma^{\mathcal{M}}A^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{y}}-g(\mu))\|^{2}}{2(\sigma^{\perp}A^{\perp})^{2}}\right)}$ (11)
For $|\operatorname{det}Dg({\bm{x}}^{*})|$, by block preservation, we can have
$Dg({\bm{x}}^{*})=\begin{bmatrix}W_{\mathcal{M}}Dg({\bm{x}}^{*})W_{\mathcal{M}}&0\\ 0&W_{\perp}Dg({\bm{x}}^{*})W_{\perp}\end{bmatrix}$ (12)
Now let $v\in{\mathcal{M}}$ be a unit vector and $\Delta$ be a scalar, then by Lipschtiz condition, we can have
$\|W_{\mathcal{M}}(g({\bm{x}}^{*}+\Delta v)-g({\bm{x}}^{*}))\|\leq B^{\mathcal{M}}\|W_{\mathcal{M}}(\Delta v)\|=B^{\mathcal{M}}|\Delta|$ (13)
With $|\Delta|\to 0$, we can have $\|W_{\mathcal{M}}Dg({\bm{x}}^{*})v\|\leq B^{\mathcal{M}}$. Since this is true for all unit vector $v\in{\mathcal{M}}$, denoting $\lambda_{i}^{\mathcal{M}}$ as the singular values of $W_{\mathcal{M}}Dg({\bm{x}}^{*})W_{\mathcal{M}}$, then for all $i\in[1,2,\dots,m]$,
$B^{\mathcal{M}}\geq\sup_{\|v\|=1}\|W_{\mathcal{M}}Dg({\bm{x}}^{*})W_{\mathcal{M}}v\|\geq s^{\mathcal{M}}_{i}$ (14)
Similarly, for singular values $\lambda_{j}^{\perp}$ of $W_{\perp}Dg({\bm{x}}^{*})W_{\perp}$, we can have
$B^{\perp}\geq\sup_{\|v\|=1}\|W_{\perp}Dg({\bm{x}}^{*})W_{\perp}v\|\geq s^{\perp}_{j}$ (15)
for all $j\in[1,2,\dots,d-m]$.
Using Equation 12,14,15, we obtain
$|\operatorname{det}Dg({\bm{x}}^{*})|=|\operatorname{det}W_{\mathcal{M}}Dg({\bm{x}}^{*})W_{\mathcal{M}}||\operatorname{det}W_{\perp}Dg({\bm{x}}^{*})W_{\perp}|=\prod_{i=1}^{m}\lambda_{i}^{\mathcal{M}}\prod_{j=1}^{d-m}\lambda_{j}^{\perp}\leq(B^{\mathcal{M}})^{m}(B^{\perp})^{d-m}$ (16)
Putting everything together, we can have
$p_{Y}({\bm{y}})\geq c(2\pi)^{-d/2}(\sigma^{\mathcal{M}}B^{\mathcal{M}})^{-m}(\sigma^{\perp}B^{\perp})^{m-d}\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{y}}-g(\mu))\|^{2}}{2(\sigma^{\mathcal{M}}A^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{y}}-g(\mu))\|^{2}}{2(\sigma^{\perp}A^{\perp})^{2}}\right)}$ (17)
To further facilitate the proof of Theorem 5.1, we further define the recursive mean as
$\mu_{T}=0,\qquad\mu_{t}=\mu_{\theta}(\mu_{t+1},t)$ (18)
and the block variance scales as
$\sigma_{T}^{\mathcal{M}}=\sigma_{T}^{\perp}=\sigma_{T}=1,\qquad(\sigma_{t-1}^{\mathcal{M}})^{2}=\sigma_{t}^{2}+(A_{t}^{\mathcal{M}})^{2}(\sigma_{t}^{\mathcal{M}})^{2},\qquad(\sigma_{t-1}^{\perp})^{2}=\sigma_{t}^{2}+(A_{t}^{\perp})^{2}(\sigma_{t}^{\perp})^{2}$ (19)
Now we are ready to prove our Theorem 5.1. Before that, we first provide the formal statement of Theorem 5.1.
Theorem B.4 (Formal statement of Theorem 5.1).
Under Assumption B.1, and let $p_{\theta}$ be a DDPM diffusion using parametrization from Equation 2 that satisfies Assumption B.2, with some constant $C$, for all ${\bm{s}}\in S$
$p_{\theta}({\bm{s}})\geq C\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{s}}-\mu_{0})\|^{2}}{2(\sigma_{0}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}{\bm{s}}\|^{2}}{2(\sigma_{0}^{\perp})^{2}}\right)}$ (20)
Proof.
Denoting $p_{\theta}({\bm{s}}_{t},t)$ as the density at time $t$ for noisy sample ${\bm{s}}_{t}$, we prove the statement by induction.
  1. 1.
    Base Case: At $t=T$, ${\bm{s}}_{T}\sim\mathcal{N}(0,I)$, therefore
    $p_{\theta}({\bm{s}}_{T},T)=(2\pi)^{-d/2}\exp{\left(-\frac{\|W_{\mathcal{M}}{\bm{s}}_{T}\|^{2}}{2}-\frac{\|W_{\perp}{\bm{s}}_{T}\|^{2}}{2}\right)}$ (21)
  2. 2.
    Inductive Hypothesis: Assume that for timestep $k$,
    $p_{\theta}({\bm{s}}_{k},k)\geq(2\pi)^{-d/2}\left(\prod_{t=k+1}^{T}(\frac{B_{t}^{\mathcal{M}}}{A_{t}^{\mathcal{M}}})^{-m}(\frac{B_{t}^{\perp}}{A_{t}^{\perp}})^{m-d}\right)(\sigma_{k}^{\mathcal{M}})^{-m}(\sigma_{k}^{\perp})^{m-d}\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{s}}-\mu_{k})\|^{2}}{2(\sigma_{k}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{s}}-\mu_{k})\|^{2}}{2(\sigma_{k}^{\perp})^{2}}\right)}$ (22)
  3. 3.
    Induction Step: For timestep k1k-1, recall ${\bm{s}}_{k-1}=\mu_{\theta}({\bm{s}}_{k},k)+\sigma_{k}\epsilon$ for $\epsilon\sim\mathcal{N}(0,I)$, we can break this down into two operations: applying $\mu_{\theta}(\cdot,k)$ and adding Gaussian white noise.
    Since $\mu_{\theta}$ satisfies Assumption B.2, let ${\bm{y}}=\mu_{\theta}({\bm{s}}_{k},k)$, we can directly apply Lemma 7 and obtain
    $p_{Y}({\bm{y}})\geq c(2\pi)^{-d/2}(\sigma_{k}^{\mathcal{M}}B_{k}^{\mathcal{M}})^{-m}(\sigma_{k}^{\perp}B_{k}^{\perp})^{m-d}\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{y}}-\mu_{\theta}(\mu_{k},k))\|^{2}}{2(\sigma_{k}^{\mathcal{M}}A_{k}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{y}}-\mu_{\theta}(\mu_{k},k))\|^{2}}{2(\sigma_{k}^{\perp}A_{k}^{\perp})^{2}}\right)}$ (23)
    where $c=\prod_{t=k+1}^{T}(\frac{B_{t}^{\mathcal{M}}}{A_{t}^{\mathcal{M}}})^{-m}(\frac{B_{t}^{\perp}}{A_{t}^{\perp}})^{m-d}$. Notice that this is equivalent to
    $p_{Y}({\bm{y}})$ $\geq c(2\pi)^{-d/2}(\sigma_{k}^{\mathcal{M}}B_{k}^{\mathcal{M}})^{-m}(\sigma_{k}^{\perp}B_{k}^{\perp})^{m-d}\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{y}}-\mu_{\theta}(\mu_{k},k))\|^{2}}{2(\sigma_{k}^{\mathcal{M}}A_{k}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{y}}-\mu_{\theta}(\mu_{k},k))\|^{2}}{2(\sigma_{k}^{\perp}A_{k}^{\perp})^{2}}\right)}$
    $=(2\pi)^{-d/2}\prod_{t=k+1}^{T}(\frac{B_{t}^{\mathcal{M}}}{A_{t}^{\mathcal{M}}})^{-m}(\frac{B_{t}^{\perp}}{A_{t}^{\perp}})^{m-d}(\sigma_{k}^{\mathcal{M}}B_{k}^{\mathcal{M}})^{-m}(\sigma_{k}^{\perp}B_{k}^{\perp})^{m-d}$
    $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{y}}-\mu_{k-1})\|^{2}}{2(\sigma_{k}^{\mathcal{M}}A_{k}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{y}}-\mu_{k-1})\|^{2}}{2(\sigma_{k}^{\perp}A_{k}^{\perp})^{2}}\right)}$
    $=(2\pi)^{-d/2}\prod_{t=k}^{T}(\frac{B_{t}^{\mathcal{M}}}{A_{t}^{\mathcal{M}}})^{-m}(\frac{B_{t}^{\perp}}{A_{t}^{\perp}})^{m-d}(\sigma_{k}^{\mathcal{M}}A_{k}^{\mathcal{M}})^{-m}(\sigma_{k}^{\perp}A_{k}^{\perp})^{m-d}$
    $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{y}}-\mu_{k-1})\|^{2}}{2(\sigma_{k}^{\mathcal{M}}A_{k}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{y}}-\mu_{k-1})\|^{2}}{2(\sigma_{k}^{\perp}A_{k}^{\perp})^{2}}\right)}$
    Adding zero-mean Gaussian white noise to $\bm{y}$} with variance $\sigma_{k}^{2}$ produces the desirable ${\bm{s}}_{k-1}$, which increase the variance in the original Gaussian bound by $\sigma_{k}^{2}$. Therefore, we can obtain
    $p_{\theta}({\bm{s}}_{k-1},k-1)$ $\geq(2\pi)^{-d/2}\left(\prod_{t=k}^{T}(\frac{B_{t}^{\mathcal{M}}}{A_{t}^{\mathcal{M}}})^{-m}(\frac{B_{t}^{\perp}}{A_{t}^{\perp}})^{m-d}\right)(\sigma_{k}^{2}+(A_{k}^{\mathcal{M}})^{2}(\sigma_{k}^{\mathcal{M}})^{2})^{-m/2}(\sigma_{k}^{2}+(A_{k}^{\perp})^{2}(\sigma_{k}^{\perp})^{2})^{(m-d)/2}$ (24)
    $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{s}}-\mu_{k-1})\|^{2}}{2(\sigma_{k}^{2}+(A_{k}^{\mathcal{M}})^{2}(\sigma_{k}^{\mathcal{M}})^{2})}-\frac{\|W_{\perp}({\bm{s}}-\mu_{k-1})\|^{2}}{2(\sigma_{k}^{2}+(A_{k}^{\mathcal{M}})^{2}(\sigma_{k}^{\mathcal{M}})^{2})}\right)}$ (25)
    $\geq(2\pi)^{-d/2}\left(\prod_{t=k}^{T}(\frac{B_{t}^{\mathcal{M}}}{A_{t}^{\mathcal{M}}})^{-m}(\frac{B_{t}^{\perp}}{A_{t}^{\perp}})^{m-d}\right)(\sigma_{k-1}^{\mathcal{M}})^{-m}(\sigma_{k-1}^{\perp})^{m-d}$ (26)
    $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{s}}-\mu_{k-1})\|^{2}}{2(\sigma_{k-1}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}({\bm{s}}-\mu_{k-1})\|^{2}}{2(\sigma_{k-1}^{\perp})^{2}}\right)}$ (27)
    Observe that since $\mu_{T}\in{\mathcal{M}}$ and $\mu_{\theta}$ is block preserving, $\mu_{t}\in{\mathcal{M}}$ for all $t$.  we can simplify the expression to be
    $p_{\theta}({\bm{s}}_{k-1},k-1)\geq(2\pi)^{-d/2}\left(\prod_{t=k}^{T}(\frac{B_{t}^{\mathcal{M}}}{A_{t}^{\mathcal{M}}})^{-m}(\frac{B_{t}^{\perp}}{A_{t}^{\perp}})^{m-d}\right)(\sigma_{k-1}^{\mathcal{M}})^{-m}(\sigma_{k-1}^{\perp})^{m-d}\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{s}}-\mu_{k-1})\|^{2}}{2(\sigma_{k-1}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}{\bm{s}}\|^{2}}{2(\sigma_{k-1}^{\perp})^{2}}\right)}$ (28)
Hence we conclude the induction.
Let $C=(2\pi)^{-d/2}\left(\prod_{t=1}^{T}(\frac{B_{t}^{\mathcal{M}}}{A_{t}^{\mathcal{M}}})^{-m}(\frac{B_{t}^{\perp}}{A_{t}^{\perp}})^{m-d}\right)(\sigma_{0}^{\mathcal{M}})^{-m}(\sigma_{0}^{\perp})^{m-d}$, we arrive at
$p_{\theta}({\bm{s}})\geq C\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{s}}-\mu_{0})\|^{2}}{2(\sigma_{0}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}{\bm{s}}\|^{2}}{2(\sigma_{0}^{\perp})^{2}}\right)}$ (29)
From this theorem above, we can make two observations: Firstly, the density assigned at a certain point depends on two factors: (1) how far it is along the manifold from the model’s noise-free center, and (2) how far it is off the manifold. Secondly, since $0<A_{t}^{\perp}\leq B_{t}^{\perp}<1$, off-manifold components incurs much harsher penalty under the DDPM model, with likelihood dropping rapidly as the off-manifold component grows. In other words, a well-trained DDPM model concentrates on in-combination states, including the OOC ones. How much density it assigns to each in-combination state is determined by how far away the state is from the model learned noise-free center.
Appendix C Experiment Details
C.1 Hardware and Platform
@@ C.4 Trajectory formulation
C.5 Maze2D
C.5.1 Extra Results
@@
C.5.2 Experiment Details
We followed the setup used in Janner et al. (2022). The hyperparameters shared for large and medium mazes are shown below in Table 1. Large maze use a planning horizon of 384 and medium maze use a planning horizon of 256. Conditioning is passed through a positional embedding layer first to map each dimension of the waypoint $(x,y,v_{x},v_{y})$ to a higher dimension of 21 and concatenate them to form a vector of size (1,214)(1,21*4). Three waypoints are then stacked together to form a matrix of size (3,214)(3,21*4) and passed into the cross-attention layer. In our experiment, directly using the waypoints as conditioning was unsuccessful.
@@ C.6.3 Dataset
C.7 Setup for Roundabout Environment
Here we describe how the Roundabout task in this paper conforms to our problem description. In this setting our base object set is $E=\{\text{car},\text{bicycle},\text{null}\}$ where null means an object is non-visible. Since the maximum number of objects in the roundabout is five and we fix the ego agent to be a car, support for the training observation is $\{(\text{car (ego agent)},\text{car},\text{car},\text{car},\text{car}),(\text{car (ego agent)},\text{bicycle},\text{bicycle},\text{bicycle},\text{bicycle})\}$ and for the testing observation is $\{(\text{car (ego agent)},\text{bicycle},\text{bicycle},\text{car},\text{car})\}$ assuming no ordering and when the state is fully observable. Since the supports for training and testing are non-overlapping under full observability, they will remain non-overlapping even when some traffic objects are out of sight, unless the ego agent is the only object present in the environment.
C.7.1 Experiment Details